OpenStudy (anonymous):

A solid is a hemisphere with 90 degree slice removed. r=8 in. volume=

Directrix (directrix):

The volume of the hemisphere is given by the following: V = (2/3)*π*r³ V = (2/3)*π*(8)³ V = ? That is the volume of the hemisphere. A 90 degree sector has been cut out. 360-90 = 270. 270/360 = 3/4 and 3/4 of the hemisphere volume remains after the cut. Take the answer you got for V and multiply it by (3/4) to get the post-cut volume of the hemisphere. @teenj12

OpenStudy (anonymous):

Thank you so much. For volume I got 1072 and when I multiplied that by 3/4 it gave me 804. Is that right?

Directrix (directrix):

That is what I got.

OpenStudy (anonymous):

Great :D. Thank you so much for the help. I really appreciate it.

Directrix (directrix):

Glad to help.